OpenStudy (anonymous):

Given: ∆BCA is a right triangle. Prove: a2 + b2 = c2 The two-column proof with missing justifications proves the Pythagorean Theorem using similar triangles.

OpenStudy (anonymous):

Statement Justification Draw an altitude from point C to AB Let = a = b = c = h = x = y y + x = a2 = cy; b2 = cx a2 + b2 = cy + b2 a2 + b2 = cy + cx a2 + b2 = c(y + x) a2 + b2 = c(c) a2 + b2 = c2 Which is NOT a justification for the proof? Addition Property of Equality Pythagorean Theorem Pieces of Right Triangles Similarity Theorem Cross Product Property